A fair six-sided die is tossed four times and the resulting sequence of numbers is
recorded.

A) What is the probability of the event E that either all four numbers are not
equal or none of them is 3 or 6?

B) What is the probability of the event E that either all four numbers are
equal or none of them is 2?

C)What is the probability of the event E that either all four numbers are not
equal and none of them is 5?
Answer #1

A) Prob(No 3 or 6 in 4 rolls or none of them are equal) = (4!)/(64) = 1/54

B) Prob( all 4 are not equal or none of them is 2) = (6*5*4*3+ 54 - 5*4*3)/ 64 = 925/1296

C) probability of the event E that either all four numbers are notequal and none of them is 5 = (5*4*3*2)/6^4 = 5/54